Jakaya received an inheritance of $70,000. She invested part at 8% and deposited the remainder in tax-free bonds at 10%. Her total annual income from the investments was $6000. Find the amount invested at 10%.
Answer by mananth(16946)

Investment I 8.00% per annum
Investment II 10.00% per annum

x+y=70000 ------------------------1
8.00%x+10.00%y= $6,000.00
Multiply by 100
8x+10y = $600,000.00 --------2

Multiply (1) by -8
we get

-8x-8y =-560000.00

Add this to (2)

2 y = $40,000.00

divide by 2

y = $20,000.00 investment at 10.00%
Balance $50,000.00 investment at 8.00%
CHECK
$50,000.00 $4,000.00
$20,000.00 $2,000.00
Total $6,000.00